OpenStudy (anonymous):

Solve: 8x − 2(x + 1) = 7x + 8

OpenStudy (anonymous):

sorry to say but im sure wrong cuz its not my option

OpenStudy (anonymous):

8x-2x-7x=8+2 -x=10 x=-10

OpenStudy (anonymous):

sorry, my bad!! that's embarrassing especially because I'm a senior D: